# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v4

@@ III-C Discussion
To the best of our knowledge, MSGT is the first GT algorithm to effectively leverage Markovian prior statistics. Unlike numerous previous approaches, MSGT utilizes initial probabilities and transition matrices without necessitating specific adjustments for new use cases. The algorithm offers the flexibility to be fine-tuned to optimize its performance in accordance with the available computational resources and the number of tests, $T$. The simple reduction of the search space in Stage 1 enables MSGT to handle challenging regimes with a small number of tests. Stage 2, and the LVA step in particular, contributes to its high success probability. Furthermore, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [32]. It is important to note that, as explained in [32], achieving optimal results is ensured only with a very large $L$, inevitably leading to complexity equivalent to MAP’s. However, as we empirically demonstrate in the following section, results equivalent to MAP’s can be achieved with reasonable complexity. Moreover, it is shown that MSGT addresses practical regimes, e.g., in COVID-19 detection [55] ($T=48$ for $(N,K)=(384,4)$), in communications [56] ($(N,K)=(10^{5},6)$), and in GT quantizer [8] ($(N,K)=(1024,16)$ [57]).
@@
Another aspect of novelty in this work is the integration of the VA into the GT problem. In the context of Markovian priors, one can think of the population’s sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps. In that case, the selection of a Viterbi decoder becomes natural, offering an optimal and efficient decoding solution. However, the most likely sequence of items does not necessarily include $K$ defective items. Particularly in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations. As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K$. To address this, we employ LVA, which produces a list of the $L$ most likely sequences, such that choosing an appropriate value for $L$ guarantees a successful recovery.
Like many previous works, MSGT assumes precise knowledge of $K$. In practical use cases, this assumption relies on using an accurate estimator for $K$ employing $\mathcal{O}\left(\log_{2}(N)\right)$ tests [58], but the estimation might be erroneous. For the sake of practical completeness, it is worth noting that modifications can be made to handle incorrect estimates of $K$, albeit with increased computational complexity. The authors of [4] suggest altering the ML estimator to consider all possible sets, without restricting the number of defective items, and show that the probability of success is almost unaffected. MSGT relies on knowing $K$ only for the MAP estimation step. Thus, a similar adaptation can be applied. In MSGT, the MAP estimator should consider all possible sets from LVA and cannot stop with the first set that explains the outcome $Y$. Hence, such modification will increase the complexity compared to [4].
@@ IV MAP Analytical Bound for GT with General Correlated Prior Statistics
In this section, we provide a sufficiency bound for the number of required pooled tests for any MAP decoder with an arbitrary underlying correlation between infected items. This bound also applies to the LVA-based MSGT algorithm, which can approach MAP decoder performance for a sufficiently large list [32]. We calculate this bound by extending the known result of ML decoding for GT problems [17], which bounds the probability of error of a ML decoder using a Gallager-type bound [53] for multiple access channel (MAC) channels [59]. Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword. While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [36, 37, 38, 39, 40, 41, 42], they do not expand naturally to GT, since GT can be thought as a special case where only “$K$” users transmit at any given point. This key difference necessitates dedicated analysis as follows in this section. In particular, to gain from the correlated prior statistics information, unlike ML decoders and bounds e.g., in [17, 14]. We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5. Finally, in Subsection IV-D, we evaluate this bound in Gilbert-Elliott (GE) model [60] and illustrate it for a practical regime in Section V.
@@ IV-A Definitions and notations
Let $\left[n\right]=\left\{1,\ldots,n\right\}$. Let $H(\cdot)$ and $I(\cdot;\cdot)$ denote Shannon’s entropy and mutual information, respectively [61]. Let $\omega\in\left[\binom{N}{K}\right]$ denote the index of a specific set of $K$ items out of the $N$ items. Throughout the paper, $\omega^{*}$ denotes the index of the set of the $K$ actual infected items. Let $W$ denote the random variable which points to the index of the infected items, and has a probability distribution of $P_{W}\left(\cdot\right):\left[\binom{N}{K}\right]\rightarrow\left[0,1\right]$. In the following, $W$ might be omitted for convenience of read. The $\omega$-th set is denoted by $S_{\omega}$. Note that $\lvert S_{\omega}\rvert=K$. The rows of $\mathbf{X}$ that correspond to the items in $S_{\omega}$ are denoted by $\mathbf{X}_{S_{\omega}}$. The output For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$. Note that $\lvert FN\left(\omega,\omega^{\prime}\right)\rvert=\lvert FP\left(\omega,\omega^{\prime}\right)\rvert$. In the following, we omit the dependency on $\omega,\omega^{*}$ and write $\mathbf{X}_{TP},\mathbf{X}_{FP},\mathbf{X}_{FN}$ instead of $\mathbf{X}_{TP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FN\left(\omega,\omega^{*}\right)}$ for notational simplicity. We say that an error of size $i$ happens if an output set of items $\omega\neq\omega^{*}$ is chosen and $\lvert FP\left(\omega,\omega^{*}\right)\rvert=i$. This event is denoted by $E_{i}$. Assuming the event $E_{i}$ happened, the index of the FN items is denoted by $j$, i.e. $j\in\left[\binom{N-K}{i}\right]$. Let $\Omega$ denote a function that converts a $j\in\left[\binom{N-K}{i}\right]$ index to a set of size $i$ of the FN items (the dependence on $\omega^{*}$ is omitted for notational simplicity). A decoder is denoted by $g$ and outputs some index $\omega=g\left(Y^{T}\right)$, where the dependency on $\mathbf{X}$ is omitted in favor of notational simplicity. These definitions are illustrated in the following example.
@@ IV-C Sufficiency GT MAP Bound
Let $(S_{2},S_{1})$ denote a partition of the defective set $S$ into disjoint sets $S_{2}$ and $S_{1}$, with cardinalities $i$ and $K-i$, respectively. This partition allows for categorizing error occurrences into distinct groups, where in group $i$, $K-i$ defective items are already identified correctly, and the primary error scenario arises from trying to detect the remaining $i$ defective items.
@@
$P(E_{i}^{j})\leq 2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
@@
Proof of Lemma 1: The proof is given in Appendix B. ∎
@@
$\frac{\partial E_{o}}{\partial\rho}|_{\rho=0}$
$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)-\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}$
$=$ $I(X_{FP};X_{TP},Y).$
And also, in Appedix C, that
$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}=H(P_{S_{2}|S_{1}}).$
This is obtained directly from [62] as also elaborated in [42, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
@@
$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
@@
For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by
@@
Hence, by applying Clain 1 in Eq. (11), we achieves the sufficiency bound on $T$ as provided in Theorem 5,
@@
(b) Lower bound for definitely defective items after DD for $N=10000$, and $K=15$.
@@ IV-D Gilbert Elliott calculations
Consider a GE model [60]. It is a Markov process with two states. In the context of GT, as also recently considered in [63], we call them infected (state 1) and not infected (state 0). The probability of transition from state 0,1 to state 1,0 is denoted by $q,s$, respectively. The probability that the first state is in state 1 is denoted by $\pi\triangleq P\left(U_{i}=1\right)$. Unless stated otherwise, we assume the system is in steady state, that is, the probability of the system to be at state 1 at any point is given by $\pi=\frac{q}{q+s}$.
@@
Let $\mathcal{G}\mathcal{E}_{N}$ denote a GE process that runs for $N$ steps. Let $\left\{U_{i}\right\}_{i=1}^{N}$ denote the states of the process during the $N$ steps. Let $\lvert\mathcal{G}\mathcal{E}_{N}\rvert\triangleq\sum_{i=1}^{N}U_{i}$. While our analysis is conditioned on the assumption that exactly $K$ items are infected, a general $\mathcal{G}\mathcal{E}_{N}$ process may result in any number of infected items between 0 and $N$. We therefore analyze and simulate a special class of GE processes; GE processes that reach the infected state exactly $K$ times after $N$ steps. Such a GE process that is guaranteed to produce exactly $K$ infected items is denoted by $\mathcal{G}\mathcal{E}_{N,K}$. Then $P_{W}\left(S_{\omega}\right)$ of $\mathcal{G}\mathcal{E}_{N,K}$ is given as follows
@@
=$\overset{(a)}{=}$ $P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\lvert\;\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)$
=$\overset{(b)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega},\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}$
=$\overset{(c)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)},$
@@
The numerator evaluates the probability that a standard GE process results in a specific pattern of states, where state one is achieved only at the states in $S_{\omega}$. Recall that GE is a Markov chain, and so $P\left(S_{\omega}\right)=P\left(U_{1}\right)\cdot\prod_{i=2}^{N}P\left(U_{i}\lvert U_{i-1}\right)$. This can be readily calculated using the definitions of $q,s,\pi$, as described at the beginning of this section
@@
We create a new Markov chain that represents $\sum_{i=1}^{n}U_{i}$ as $n$ advances from $1$ to $N$. This Markov chain keeps track of how many times the GE process was in state 1. Hence this Markov chain has $2\left(N+1\right)$ states. State $n,n^{\prime}\in\left\{0,\ldots,N\right\}$ denotes a state where the GE process was in the infected state (state 1) $n$ times, and the current state is 0,1, respectively. An example of this Markov chain is shown in Fig. 6. For convenience, the states $0^{\prime},N$ are also listed, even though they cannot be achieved. This Markov chain’s states can be arranged in a vector, where the $j$-th state represents $\sum_{i=1}^{N}U_{i}=\left\lfloor j/2\right\rfloor$, and even/odd $j$ represents the current state is not/is 00. For instance, in the example of Fig. 6, the states $6/7$ represent states 2/$2/2^{\prime}$ in the figure. The initial state vector is given as by $\left(1-\pi,0,0,\pi,0,\ldots,0\right)$. The transition matrix is given as follows
@@
Figure 5: Success probability of MSGT, MAP, ML and DD over 1000 iterations. A comparison to ML and MAP is not presented in (b) and (c), as they are not feasible for populations of those sizes due to the computational complexity burden.
@@
To conclude the calculation, the probability to have exactly $K$ infected items after $N$ steps can be obtained by summing items $12K+$ 1 and $22K+$ 2 of the vector in (13).
Figure 6: Markov chain for the number of infected items of a GE model, $N=4$.
In the next section, we evaluate the proposed sufficient bound given in Theorem 5 for MAP decoder in the GE setting considered herein.

### v5

@@ III-C Discussion
To the best of our knowledge, 2SDGT is the first GT algorithm to effectively leverage Markovian prior statistics. Unlike numerous previous approaches, 2SDGT uses the same decoding flow for any prior model, without requiring problem-specific redesign. In contrast, prior non-adaptive GT algorithms that incorporate structural assumptions—such as graph-constrained GT [17], model-based and graph-based GT [39], and contact-based GT [34]—must often adapt their test design or graph structure when the population model changes. 2SDGT instead relies only on the given initial distribution and transition matrices, while the algorithmic pipeline remains unchanged. The algorithm offers the flexibility to be fine-tuned to optimize its performance in accordance with the available computational resources and the number of tests, $T$. The simple reduction of the search space in Stage 1 enables 2SDGT to handle challenging regimes with a small number of tests. Stage 2, particularly the LVA step, contributes to its high success probability. Additionally, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [56]. It is important to note that, as explained in [56], achieving optimal results is ensured only with a very large $L$, inevitably leading to complexity equivalent to MAP’s. However, as we empirically demonstrate in the following section, results equivalent to MAP’s can be achieved with reasonable complexity. Moreover, it is shown that 2SDGT addresses practical regimes, e.g., in COVID-19 detection [57] ($T=48$ for $(N,K)=(384,4)$), in communications [45] ($(N,K)=(10^{5},6)$), and in GT quantizer [20] ($(N,K)=(1024,16)$ [36]).
@@
Another aspect of novelty in this research is the integration of the Viterbi Algorithm into the GT problem. In the context of Markovian priors, one can think of the population sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps. In that case, the selection of a Viterbi decoder becomes natural, offering an optimal and efficient decoding solution. However, the most likely sequence of items does not necessarily include $K_{\max}$ defective items. Particularly, in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations. As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K_{\max}$. To address this, we employ LVA, which produces a list of the $L$ most likely sequences, such that choosing an appropriate value for $L$ guarantees a successful recovery.
Unlike works that assume a deterministic or perfectly known $K$, the model adopted in this work allows the realized number of defectives $K$ to be random and unknown to the decoder. To accommodate this uncertainty without redesigning the algorithmic pipeline, 2SDGT operates with a design interval $(K_{\min},K_{\max})$, which is used consistently for test design, Stage 2 filtering, and candidate enumeration.
Specifically, in Stage 2, the decoder retains only trajectories whose estimated support size satisfies $K_{\min}\leq|\mathbf{V}^{(l)}|\leq{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$, and enumerates candidate subsets of all cardinalities $k\in[K_{\min},K_{\max}]$ within these supports. This modification eliminates the need to know the exact realization of $K$ while preserving the counting-based guarantees and complexity bounds derived in this Section. The overall recovery error probability under random $K$ is controlled by conditioning on the typical event $\mathcal{E}_{\mathrm{typ}}\triangleq\{K_{\min}\leq K\leq K_{\max}\}$, and adding the probability of its complement. In particular,
$P_{e}\;\leq\;P_{e}\big|\mathcal{E}_{\mathrm{typ}}\;+\;\Pr(K<K_{\min})\;+\;\Pr(K>K_{\max}).$
When $\{U_{i}\}_{i=1}^{N}$ are generated by an ergodic finite-state Markov source (or, more generally, a strongly mixing process) [24, 29], the random variable $K=\sum_{i=1}^{N}U_{i}$ satisfies concentration bounds analogous to Chernoff inequalities [40, 49]. That is, the concentration bounds invoked below hold under the additional assumption that the $\tau$-order Markov prior is ergodic (irreducible and aperiodic), or more generally strongly mixing; this assumption is standard in correlated-source models and is used solely to control the tail behavior of $K$. In particular, letting $\mu=\mathbb{E}[K]$, there exist constants $c_{1},c_{2}>0$ such that for any $0<\varepsilon\leq 1$, $\Pr\!\left(K>(1+\varepsilon)\mu\right)\leq c_{1}\exp\!\left(-c_{2}\,\varepsilon^{2}\,\mu\right)$, and $\Pr\!\left(K<(1-\varepsilon)\mu\right)\leq c_{1}\exp\!\left(-c_{2}\,\varepsilon^{2}\,\mu\right)$. Therefore, choosing $K_{\min}=\lfloor(1-\varepsilon)\mu\rfloor$, and $K_{\max}=\lceil(1+\varepsilon)\mu\rceil$, one can obtain that both tail probabilities decay exponentially in $\mu$. As a result, for sufficiently sparse regimes ($\mu=\mathcal{O}(N^{\theta})$ with $\theta<1$), the probability that $K$ falls outside the design interval $(K_{\min},K_{\max})$ is negligible, while 2SDGT retains its correctness and feasible computational complexity guarantees. If the above ergodicity or strong mixing assumptions do not hold, alternative concentration bounds or tail estimates for dependent random variables may be employed to control $\Pr(K<K_{\min})$ and $\Pr(K>K_{\max})$, albeit potentially with weaker decay rates.
@@ IV MAP Analytical Bound for GT with General Correlated Prior Statistics
In this section, we revisit the classical information-theoretic sufficiency condition for MAP decoding in group testing and adapt it to the setting considered here, which involves general correlated prior statistics and a constrained typical pattern in which only $K_{\min}\leq K\leq K_{\max}$ out of $N$ items are defective. This bound also applies to the LVA-based 2SDGT algorithm, which can approach MAP decoder performance for a sufficiently large list [56]. We calculate this bound by extending the known result of ML decoding for GT problems [8], which bounds the probability of error of a ML decoder using a Gallager-type bound [29] for multiple access channel (MAC) channels [30]. Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword. While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [58, 69, 70, 12, 11, 51, 52], they do not expand naturally to GT, since GT can be thought as a special case where only “KK” users transmit simultaneously at any given point. This key difference necessitates dedicated analysis as follows in this section. In particular, to gain from the correlated prior statistical information, unlike ML decoders and bounds e.g., in [8, 53, 4]. That is, unlike prior works that analyze ML decoding under a uniform prior over defective sets of fixed cardinality, our analysis considers MAP decoding with arbitrary, possibly correlated priors, requiring explicit weighting of competing defective sets according to their prior probabilities. We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5. In Subsection IV-D, we evaluate this bound in Gilbert-Elliott model [33] and illustrate it for a practical regime in Section V. Throughout those subsection, for essay of notation and to follow the existing GT literature, we first analyze the performance of a MAP decoder conditioned on a fixed realization $K=k$. Accordingly, all quantities below (including the index set $\left[\binom{N}{K}\right$], the random variable $W$, and its distribution $P_{W}$) are interpreted under the conditional prior $P(\cdot\mid K=k)$, and the resulting bound applies to the conditional error probability $P_{e}(k)\triangleq\Pr\!\big(\widehat{\mathcal{K}}\neq\mathcal{K}\,\big|\,K=k\big)$. Finally, in Subsection IV-E, we extend the analysis to support the random-$K$ model (see Section II) with tail probability (See Section III-C).
@@ IV-A Definitions and notations
Let $\left[n\right]=\left\{1,\ldots,n\right\}$. Let $H(\cdot)$ and $I(\cdot;\cdot)$ denote Shannon’s entropy and mutual information, respectively [23]. Let $\omega\in\left[\binom{N}{K}\right$] denote the index of a specific set of $K$ items out of the $N$ items. Throughout the paper, $\omega^{*}$ denotes the index of the set of the $K$ actual infected items. Let $W$ denote the random variable which points to the index of the infected items, and has a probability distribution of $P_{W}\left(\cdot\right):\left[\binom{N}{K}\right]\rightarrow\left[0,1\right]$. In the following, $W$ might be omitted for convenience of read. The $\omega$-th set is denoted by $S_{\omega}$ [^4] Note that there is a one-to-one mapping from $\omega$ to $S_{\omega}$, so in the MAP bound analysis herein, we use both $P_{W}\left(S_{\omega}\right)$ and $P_{W}\left(\omega\right)$ interchangeably.. Note that $\lvert S_{\omega}\rvert=K$. The rows of $\mathbf{X}$ that correspond to the items in $S_{\omega}$ are denoted by $\mathbf{X}_{S_{\omega}}$. For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$. Note that $\lvert FN\left(\omega,\omega^{\prime}\right)\rvert=\lvert FP\left(\omega,\omega^{\prime}\right)\rvert$. In the following, we omit the dependency on $\omega,\omega^{*}$ and write $\mathbf{X}_{TP},\mathbf{X}_{FP},\mathbf{X}_{FN}$ instead of $\mathbf{X}_{TP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FN\left(\omega,\omega^{*}\right)}$ for notational simplicity. We say that an error of size $i$ happens if an output set of items $\omega\neq\omega^{*}$ is chosen and $\lvert FP\left(\omega,\omega^{*}\right)\rvert=i$. This event is denoted by $E_{i}$. Assuming the event $E_{i}$ happened, the index of the FN items is denoted by $j$, i.e. $j\in\left[\binom{N-K}{i}\right$]. Let $\Omega$ denote a function that converts a $j\in\left[\binom{N-K}{i}\right$] index to a set of size $i$ of the FN items (the dependence on $\omega^{*}$ is omitted for notational simplicity). A decoder is denoted by $g$ and outputs some index $\omega=g\left(Y^{T}\right)$, where the dependency on $\mathbf{X}$ is omitted in favor of notational simplicity. These definitions are illustrated in the following example.
@@ IV-C Sufficiency GT MAP Bound
Let $(S_{2},S_{1})$ denote a partition of the defective set $S$ into disjoint sets $S_{2}$ and $S_{1}$, with cardinalities $i$ and KiK-i, respectively. This partition allows for categorizing error occurrences into distinct groups, where in group $i$, KiK-i defective items are already identified correctly, and the primary error scenario arises from trying to detect the remaining $i$ defective items.
@@
Remark 5.
While Theorem 5 is stated for a MAP decoder conditioned on a fixed realization $K=k$, it forms the analytical foundation for the random-$K$ model adopted throughout the paper. As discussed in Section III-C and presented in Section IV-E, the overall error probability under random $K$ is controlled by conditioning on a typical sparsity interval and adding exponentially decaying tail probabilities. Finally, since the proposed 2SDGT algorithm approaches MAP performance for sufficiently large list sizes, this bound also characterizes the fundamental test-complexity regime in which 2SDGT operates effectively.
@@
$P(E_{i}^{j})\leq 2^{-T\Big(E_{0}(\rho)-{\color[rgb]{0,0,0}\left(1/\binom{K}{i}\right)E_{s,j}(\rho,P_{W})}/{T}\Big)},$
@@
Proof of Lemma 1: The proof is given in Appendix C. ∎
@@
$\hskip-24.18501pt\frac{\partial E_{o}}{\partial\rho}|_{\rho=0}$
$\hskip-19.91684pt=$ $\hskip-14.22636pt\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)-\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
$\hskip-19.91684pt=$ $\hskip-14.22636pt\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}$
$\hskip-19.91684pt=$ $\hskip-14.22636ptI(X_{FP};X_{TP},Y).$
And also, in Appedix D, that
$\hskip-24.18501pt\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}=H(P_{S_{2}|S_{1}}).$
This is obtained directly from [59] as also elaborated in [52, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
@@
$\hskip-14.22636ptT\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}+T\rho\Bigg(I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg)$
@@
Under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by
@@
Claim 1 relies exclusively on the independence of the Bernoulli test design used to construct the testing matrix $X$. No independence is assumed for the defective-item prior. In particular, the result holds for arbitrary (possibly correlated) priors on $\mathbf{U}$, including the trellis-based model considered in this paper. All prior dependence is accounted for in the MAP decoding step through the enumeration and weighting of competing defective supports, while Claim 1 characterizes only the conditional likelihood of the test outcomes given a fixed support. Hence, by applying Claim 1 in Eq. (14), we achieves the sufficiency bound on $T$ as provided in Theorem 5,
@@
(b) Lower bound for definitely defective items after $D$ for $N=10000$, and $K=15$.
@@ IV-D Gilbert Elliott calculations
Consider a Gilbert-Elliott model [33]. It is a Markov process with two states. In the context of GT, as also recently considered in [50], we call them infected (state 1) and not infected (state 0). The probability of transition from state 0,1 to state 1,0 is denoted by $q,s$, respectively. The probability that the first state is in state 1 is denoted by $\pi\triangleq P\left(U_{i}=1\right)$. Unless stated otherwise, we assume the system is in steady state, that is, the probability of the system to be at state 1 at any point is given by $\pi=\frac{q}{q+s}$.
@@
Let $\mathcal{G}\mathcal{E}_{N}$ denote a Gilbert-Elliott process that runs for $N$ steps. Let $\left\{U_{i}\right\}_{i=1}^{N}$ denote the states of the process during the $N$ steps. Let $\lvert\mathcal{G}\mathcal{E}_{N}\rvert\triangleq\sum_{i=1}^{N}U_{i}$. While our analysis is conditioned on the assumption that exactly $K$ items are infected, a general $\mathcal{G}\mathcal{E}_{N}$ process may result in any number of infected items between 0 and $N$. We therefore analyze and simulate a special class of Gilbert-Elliott processes; Gilbert-Elliott processes that reach the infected state exactly $K$ times after $N$ steps. Such a Gilbert-Elliott process that is guaranteed to produce exactly $K$ infected items is denoted by $\mathcal{G}\mathcal{E}_{N,K}$. Then $P_{W}\left(S_{\omega}\right)$ of $\mathcal{G}\mathcal{E}_{N,K}$ is given as follows
@@
$\overset{(a)}{=}$ $P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\lvert\;\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)$
$\overset{(b)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega},\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}$
$\overset{(c)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)},$
@@
The numerator evaluates the probability that a standard Gilbert-Elliott process results in a specific pattern of states, where state one is achieved only at the states in $S_{\omega}$. Recall that Gilbert-Elliott is a Markov chain, and so $P\left(S_{\omega}\right)=P\left(U_{1}\right)\cdot\prod_{i=2}^{N}P\left(U_{i}\lvert U_{i-1}\right)$. This can be readily calculated using the definitions of $q,s,\pi$, as described at the beginning of this section
@@
We create a new Markov chain that represents $\sum_{i=1}^{n}U_{i}$ as $n$ advances from $1$ to $N$. This Markov chain keeps track of how many times the Gilbert-Elliott process was in state 1. Hence this Markov chain has $2\left(N+1\right)$ states. State $n,n^{\prime}\in\left\{0,\ldots,N\right\}$ denotes a state where the Gilbert-Elliott process was in the infected state (state 1) $n$ times, and the current state is 0,1, respectively. An example of this Markov chain is shown in Fig. 7. For convenience, the states $0^{\prime},N$ are also listed, even though they cannot be achieved. This Markov chain’s states can be arranged in a vector, where the $j$-th state represents $\sum_{i=1}^{N}U_{i}=\left\lfloor j/2\right\rfloor$, and even/odd $j$ represents the current state is not/is 0. For instance, in the example of Fig. 7, the states $6/7$ represent states $2/2^{\prime}$ in the figure. The initial state vector is given as by $\left(1-\pi,0,0,\pi,0,\ldots,0\right)$. The transition matrix is given as follows
@@
Figure 6: Success probability of 2SDGT, MAP, ML, SCOMP, and DD over 1000 iterations. A comparison to ML and MAP is not presented in (b) and (c), as they are not feasible for populations of those sizes due to the computational complexity burden.
@@
To conclude the calculation, the probability to have exactly $K$ infected items after $N$ steps can be obtained by summing items $2K+1$ and $2K+2$ of the vector in (16).
Figure 7: Markov chain for the number of infected items of a Gilbert-Elliott model, $N=4$.
In the next section, we evaluate the proposed sufficient bound given in Theorem 5 for MAP decoder in the Gilbert-Elliott setting considered herein.
IV-E Extension to random-$K$ with tail probability
Here we extend the analysis of the sufficient bound for fixed defectives $K=k$ to the model given in Section II, where the defectives $K$ is a random variable. The analyses in the subsection before are conditional on the realization $K=k$ and yield a conditional MAP error bound $\Pr(\widehat{\mathcal{K}}\neq\mathcal{K}\mid K=k)$. By the law of total probability, the overall error probability satisfies
$\Pr(\widehat{\mathcal{K}}\neq\mathcal{K})=\sum_{k=K_{\min}}^{K_{\max}}\Pr(K=k)\Pr(\widehat{\mathcal{K}}\neq\mathcal{K}\mid K=k)+\Pr\!\big(K\notin[K_{\min},K_{\max}]\big).$
Consequently,
$\Pr(\widehat{\mathcal{K}}\neq\mathcal{K})\leq\max_{K_{\min}\leq k\leq K_{\max}}\Pr(\widehat{\mathcal{K}}\neq\mathcal{K}\mid K=k)+\Pr\!\big(K\notin[K_{\min},K_{\max}]\big).$
Hence, if the condition of Theorem 5 holds uniformly for all $k\in[K_{\min},K_{\max}]$ and the tail probability $\Pr(K\notin[K_{\min},K_{\max}])$ is negligible, then reliable recovery is guaranteed.